OpenStudy (anonymous):

What is the factored form of the expression? k2 – 9h2 A. (k + 3h)(k + 3h) B. h2(k + 3)(k – 3) C. (k + 3h)(k – 3h) D. (k – 3h2)(k + 3)

OpenStudy (anonymous):

this is difference of 2 squares a^2 - b^ = (a - b)( a + b) - this should help

OpenStudy (anonymous):

it is in the formula of a^2 -b^2

OpenStudy (anonymous):

a^2 - b^2

OpenStudy (anonymous):

so which one

OpenStudy (anonymous):

compare the 2 below a^2 - b^2 = (a + b)(a - b) k^2 - 9h^2 =

OpenStudy (anonymous):

so sorry its C

OpenStudy (anonymous):

sheg?

OpenStudy (anonymous):

@joeywhite sorry buddy

OpenStudy (anonymous):

lol - its ok